# Hopf orbits and the first ECH capacity

### v2

@@ 2. The Hopf capacity
Let $X\subset\mathbb{R}^{4}$ be a star-shaped domain, and let $\lambda=\lambda_{0}|_{\partial X}$ be the induced contact for on $\partial X$. Then $d\lambda$ restricts to a symplectic structure on the contact structure $\xi=\operatorname{Ker}(\lambda)$. We denote by $\phi^{t}:\partial X\to\partial X$ the flow of the Reeb vector field $R_{\lambda}$. The derivative of the flow restricts to a symplectic linear map
@@
We say that the domain $X$ (or the contact form on $\partial X$) is nondegenerate (resp. nondegenerate up to action $C$) if all Reeb orbits $\gamma\in\mathcal{P}(\partial X)$ (resp. all Reeb orbits with action less than $C$) are nondegenerate. Finally, we say that the domain $X$ (or the contact form on $\partial X$) is dynamically convex up to action $C$ if ${\rm CZ}(\gamma)\geq 3$ for all $\gamma\in\mathcal{P}(\partial X)$ with $\mathcal{A}(\gamma)<C$.
@@
Let $\Omega_{1}$ and $\Omega_{2}$ be star-shaped domains in $\mathbb{R}^{4}$. Let $C>\mathcal{A}_{\rm Hopf}(\Omega_{2})$ and assume that $\Omega_{1},\Omega_{2}$ are, up to action $C$, nondegenerate and dynamically convex. Suppose that there exists a symplectic embedding $\varphi:\Omega_{1}\hookrightarrow\mathrm{int}(\Omega_{2})$. Then there exists a Hopf orbit $\gamma\in\mathcal{P}_{\rm Hopf}(\Omega_{1})$ such that ${\rm CZ}(\gamma)=3$ and $\mathcal{A}(\gamma)<\mathcal{A}_{\rm Hopf}(\Omega_{2})$. In particular, $\mathcal{A}_{\rm Hopf}(\Omega_{1})<\mathcal{A}_{\rm Hopf}(\Omega_{2})$.
@@
We proceed in five steps.
@@
We now set up some notation.
If $g:S^{3}\to\mathbb{R}^{>0}$ is a smooth function, define a star-shaped domain
$\Omega_{g}=\{z\in\mathbb{R}^{4}\mid|z|\leq g(z/|z|)\}.$
Every star-shaped domain in $\mathbb{R}^{4}$ can be uniquely described this way. There is a diffeomorphism $S^{3}\longrightarrow\partial\Omega_{g}$ sending $z\mapsto g(z)z$. This diffeomorphism pulls back the contact form on ${\lambda_{0}}|_{\partial X_{g}}$ to $g^{2}({\lambda_{0}}|_{S^{3}})$.
@@
We now show that if $X\subset\mathbb{R}^{4}$ is a dynamically convex domain which is nondegenerate up to action $C>\mathcal{A}_{\rm Hopf}(X)$, then
(2.3) $c_{\rm Hopf}(X)=\mathcal{A}_{\rm Hopf}(X).$
To prove (2.3), first note that it follows from (1.8) that $c_{\rm Hopf}(X)\geq\mathcal{A}_{\rm Hopf}(X)$. To prove the reverse inequality, suppose that $\Omega$ is dynamically convex and that there exists a symplectic embedding $\Omega\hookrightarrow X$. We need to show that
To do so, let $\varepsilon>0$ be given. Let $\lambda$ denote the contact form on $\partial\Omega$. By a standard genericity argument, there is a function
$f:\partial\Omega\longrightarrow(1,1+\varepsilon)$
such that the contact form $f^{2}\lambda$ is nondenegerate. By a compactness argument (compare Step 5 below), we can choose $f$ so that $f^{2}\lambda$ is also dynamically convex up to action $C$. By Step 2 we can write $\Omega=\Omega_{g}$ for some $g:S^{3}\to\mathbb{R}^{>0}$. We then have symplectic embeddings
$\Omega\hookrightarrow\Omega_{fg}\hookrightarrow(1+\varepsilon)X.$
Applying Lemma 2.1 twice and equation (2.2), we obtain
$\mathcal{A}_{\rm Hopf}(\Omega)<\mathcal{A}_{\rm Hopf}(\Omega_{fg})<\mathcal{A}_{\rm Hopf}((1+\varepsilon)X)=(1+\varepsilon)^{2}\mathcal{A}_{\rm Hopf}(X).$
Since $\varepsilon>0$ was arbitrary, the inequality (2.4) follows.
Step 4
@@
There are two simple Reeb orbits in $\partial E(a,b)$, given by the circles $(z_{2}=0)$ and $(z_{1}=0)$. These Reeb orbits have actions $a$ and $b$ respectively, and they are both Hopf orbits. If $a/b$ is rational then there are additional simple Reeb orbits, but these do not have minimal action. Using Step 3, we conclude that
@@
To prove that $c_{\operatorname{Hopf}}(Z(R))=R$, let $\Omega\subset\mathbb{R}^{4}$ be a dynamically convex domain and suppose that there exists a symplectic embedding $\varphi:\Omega\hookrightarrow Z(R)$. We need to show that $\mathcal{A}_{\operatorname{Hopf}}(\Omega)\leq R$. Given $\varepsilon>0$, there exists $C>>0$ such that $\varphi(\Omega)\subset E((1+\varepsilon)R,C)$. By Step 3, Monotonicity, and equation (2.5), it follows that
@@
Step 5
To complete the proof of Theorem 1.12, we need to show that if $X\subset\mathbb{R}^{4}$ is a dynamically convex domain, then there exists $\gamma\in\mathcal{P}_{\rm Hopf}(\partial X)$ with ${\rm CZ}(\gamma)=3$ such that $\mathcal{A}(\gamma)=\mathcal{A}_{\rm Hopf}(X)$.
Let $\{\gamma_{n}\}_{n=1,\ldots}$ be a sequence of Hopf orbits in $\mathcal{P}_{\rm Hopf}(\partial X)$ with $\mathcal{A}(\gamma_{n})<e\mathcal{A}_{\operatorname{Hopf}}(X)$ and $\lim_{n\to\infty}\mathcal{A}(\gamma_{n})=\mathcal{A}_{\operatorname{Hopf}}(X)$. We remark that by a compactness argument and the Arzeli-Ascoli theorem, one can pass to a subsequence so that $\gamma_{n}$ converges to a Reeb orbit with action $\mathcal{A}_{\operatorname{Hopf}}(X)$. However this not enough to prove the theorem, because the limiting Reeb orbit might not be simple or have Conley-Zehnder index 3. We will instead need to take a limit of a sequence of Reeb orbits of different contact forms.
Let $\lambda$ denote the contact form on $\partial X$. By Step 2, replacing $X$ with a different star-shaped domain is equivalent to multiplying $\lambda$ by a positive function. We can construct a sequence of functions $f_{n}:\partial X\to(0,1)$ with $f_{n}\to 0$ in $C^{\infty}$ such that:
@@
  • The Reeb vector field of $e^{f_{n}}\lambda$ is tangent to $\gamma_{n}$, so that the loop $\gamma_{n}$, suitably reparametrized, defines a Reeb orbit $\widetilde{\gamma}_{n}\in\mathcal{P}_{\operatorname{Hopf}}(e^{f_{n}}\lambda)$.
Note that
@@
(2.7) $\lim_{n\to\infty}\mathcal{A}(\widetilde{\gamma}_{n})=\mathcal{A}_{\operatorname{Hopf}}(X).$
Now for each $n$, choose a constant $\epsilon_{n}\in(0,\min(f_{n}))$. The contact form $e^{f_{n}-\epsilon_{n}}\lambda$ is nondegenerate and dynamically convex up to action $e^{2}\mathcal{A}_{\operatorname{Hopf}}(X)$, since it is a constant scaling of $e^{f_{n}}\lambda$. By Lemma 2.1 and (2.6), there is a Hopf orbit $\gamma_{n}^{\prime}\in\mathcal{P}_{\operatorname{Hopf}}(e^{f_{n}-\epsilon_{n}}\lambda)$ such that
(2.8) $\mathcal{A}(\gamma_{n}^{\prime})<\mathcal{A}(\widetilde{\gamma}_{n})$
and $\operatorname{CZ}(\gamma_{n}^{\prime})=3$.
Now by a compactness argument using the Arzela-Ascoli theorem, we can pass to a subsequence so that $\gamma_{n}^{\prime}$ converges to a Reeb orbit $\gamma\in\mathcal{P}(\partial X)$. By (2.7) and (2.8), we have $\mathcal{A}(\gamma)\leq\mathcal{A}_{\operatorname{Hopf}}(X)$. So to complete the proof, we just need to show that $\gamma$ is a Hopf orbit and $\operatorname{CZ}(\gamma)=3$.
By dynamical convexity and the lower semicontinuity of the Conley-Zehnder index, we have ${\rm CZ}(\gamma)=3$. If $\gamma$ is a simple Reeb orbit, then $\gamma$ is a Hopf orbit and we are done. If $\gamma$ is not simple, then $\gamma$ is a multiple cover of some Reeb orbit $\gamma^{\prime}$, which by dynamical convexity has rotation number $\rho(\gamma^{\prime})>1$. It follows that the rotation number $\rho(\gamma)>2$, so ${\rm CZ}(\gamma)\geq 5$, which is a contradiction. ∎
We now prepare for the proof of Lemma 2.1, which occupies the rest of §2.
@@ 2.2. Reeb orbits, asymptotic operators, and Conley-Zehnder indices
[…]
@@
(2.9) ${\rm CZ}^{\delta}_{\tau}(\gamma)=2\alpha^{<\delta}_{\tau}(\gamma)+p^{\delta}(\gamma).$
We note that ${\rm CZ}_{\tau}^{\delta}$ is a lower semi-continuous function of the asymptotic operator $L_{\gamma,J}$ with respect to the $C^{\infty}$ norm. We define the Conley-Zehnder index ${\rm CZ}_{\tau}(\gamma)={\rm CZ}_{\tau}^{0}(\gamma)$, i.e. by taking $\delta=0$ in (2.9).
@@
${\rm CZ}(\gamma)={\rm CZ}_{\tau}^{0}(\gamma).$

### v3

@@ 2. The Hopf capacity
Let $X\subset\mathbb{R}^{4}$ be a star-shaped domain, and let $\lambda=\lambda_{0}|_{\partial X}$ be the induced contact form on $\partial X$. Then $d\lambda$ restricts to a symplectic structure on the contact structure $\xi=\operatorname{Ker}(\lambda)$. We denote by $\phi^{t}:\partial X\to\partial X$ the flow of the Reeb vector field $R_{\lambda}$. The derivative of the flow restricts to a symplectic linear map
@@
We say that the domain $X$ (or the contact form on $\partial X$) is nondegenerate (resp. nondegenerate up to action $C$) if all Reeb orbits $\gamma\in\mathcal{P}(\partial X)$ (resp. all Reeb orbits with action less than $C$) are nondegenerate. Finally, we say that the domain $X$ (or the contact form on $\partial X$) is dynamically convex up to action $C$ if $\operatorname{CZ}(\gamma)\geq 3$ for all $\gamma\in\mathcal{P}(\partial X)$ with $\mathcal{A}(\gamma)\leq C$.
@@
Let $\Omega_{1}$ and $\Omega_{2}$ be star-shaped domains in $\mathbb{R}^{4}$. Let $C>\mathcal{A}_{\rm Hopf}(\Omega_{2})$ and assume that:
  • $\Omega_{1}$ is nondegenerate up to action $C$.
  • $\Omega_{1}$, $\Omega_{2}$ are dynamically convex up to action $C$.
Suppose that there exists a symplectic embedding $\varphi:\Omega_{1}\to\Omega_{2}$. Then there exists a Hopf orbit $\gamma\in\mathcal{P}_{\rm Hopf}(\Omega_{1})$ such that $\operatorname{CZ}(\gamma)=3$ and $\mathcal{A}(\gamma)\leq\mathcal{A}_{\rm Hopf}(\Omega_{2})$. In particular, $\mathcal{A}_{\rm Hopf}(\Omega_{1})\leq\mathcal{A}_{\rm Hopf}(\Omega_{2})$. If, in addition, $\Omega_{2}$ is nondegenerate up to action $C$ and $\varphi(\Omega_{1})\subset{\rm int}(\Omega_{2})$ then $\mathcal{A}(\gamma)<\mathcal{A}_{\rm Hopf}(\Omega_{2})$.
@@
We proceed in four steps.
@@
We now show that if $X\subset\mathbb{R}^{4}$ is any dynamically convex domain then
(2.3) $c_{\rm Hopf}(X)=\mathcal{A}_{\rm Hopf}(X).$
To prove (2.3), first we need an auxiliary lemma.
Lemma 2.2.
Let $\Omega\subset\mathbb{R}^{4}$ be a star-shaped domain that is dynamically convex up to $C>\mathcal{A}_{\rm Hopf}(\Omega)$. Let $\Omega_{n}\subset\mathbb{R}^{4}$ be star-shaped domains satisfying $\partial\Omega_{n}\to\partial\Omega$ in $C^{\infty}$ as hypersurfaces. Let $\gamma_{n}\in\mathcal{P}_{\rm Hopf}(\partial\Omega_{n})$ satisfy $\sup_{n}\mathcal{A}(\gamma_{n})\leq C$ and ${\rm CZ}(\gamma_{n})=3\ \forall n$. There exists $\gamma\in\mathcal{P}_{\rm Hopf}(\partial\Omega)$ satisfying $\mathcal{A}(\gamma)=\liminf_{n\to\infty}\mathcal{A}(\gamma_{n})$ and ${\rm CZ}(\gamma)=3$.
It is important to note that the above statement makes no nondegeneracy assumptions.
Proof of Lemma 2.2.
The statement can be rewritten in terms of contact forms on $S^{3}$ defining its standard contact structure $\xi_{0}=\ker\lambda_{0}|_{S^{3}}$. Let $\lambda$, $\lambda_{1},\lambda_{2},\dots$ be contact forms on $(S^{3},\xi_{0})$. Assume that $\lambda$ is dynamically convex up to action $C$ and that $\lambda_{n}\to\lambda$ in $C^{\infty}$. Let $\gamma_{n}$ be periodic $\lambda_{n}$-Reeb orbits which are Hopf orbits, i.e. each $\gamma_{n}$ is simple, unknotted and has self-linking number 1-1. Assume that ${\rm CZ}(\gamma_{n})=3\ \forall n$ and that $\sup_{n}\int_{\gamma_{n}}\lambda_{n}\leq C$. We need to show that there exists a simple periodic $\lambda$-Reeb orbit $\gamma$ which is unknotted, has self-linking number 1-1, and $\int_{\gamma}\lambda=\liminf_{n\to\infty}\int_{\gamma_{n}}\lambda_{n}$. In fact, since $\int_{\gamma_{n}}\lambda_{n}$ is the period of $\gamma_{n}$ as a $\lambda_{n}$-Reeb orbit, we can apply the Arzelà-Ascoli Theorem and assume, without loss of generality and up to selection of a subsequence, that for some periodic $\lambda$-Reeb orbit $\gamma$ we have $\gamma_{n}\to\gamma$ in $C^{\infty}$. The $\gamma_{n}$ are simple by assumption but, in principle, it might not be the case that $\gamma$ is simple. If $\gamma$ is not simple then, by the dynamical convexity of $\lambda$ up to action $C$ and the fact that $\int_{\gamma}\lambda\leq C$, we have ${\rm CZ}(\gamma)\geq 5$. However, the lower-semicontinuity of the Conley-Zehnder index and the dynamical convexity of $\Omega$ up to action $C$ together imply that $3\leq{\rm CZ}(\gamma)\leq\liminf_{n\to\infty}{\rm CZ}(\gamma_{n})=3$. This shows that $\gamma$ is a simple periodic orbit with ${\rm CZ}(\gamma)=3$, and that $\gamma$ is transversely isotopic to $\gamma_{n}$ for $n$ large enough. Hence, $\gamma$ is unknotted with self-linking number 1-1. ∎
Suppose that $\Omega$ is dynamically convex and that there exists a symplectic embedding $\Omega\hookrightarrow X$. We will now show that
Let $\lambda$ denote the contact form on $\partial\Omega$. Consider $\Omega_{n}$ nondegenerate star-shaped domains satisfying $\Omega_{n}\subset{\rm int}(\Omega)$, and $\Omega_{n}\to\Omega$ in $C^{\infty}$. We then have symplectic embeddings $\Omega_{n}\hookrightarrow\mathrm{int}(X)$. For every $C>0$ there exists $n_{C}\in\mathbb{N}$ such that if $n\geq n_{C}$ then $\Omega_{n}$ is dynamically convex up to action $C$. This is a simple consequence of the Arzelà-Ascoli Theorem and the fact that the Conley-Zehnder index is lower-semicontinuous. By Lemma 2.1, we find $\gamma_{n}\in\mathcal{P}_{\rm Hopf}(\partial\Omega_{n})$ with ${\rm CZ}(\gamma_{n})=3$ and $\mathcal{A}(\gamma_{n})\leq\mathcal{A}_{\rm Hopf}(X)$. By an application of Lemma 2.2, there exists $\gamma\in\mathcal{P}_{\rm Hopf}(\partial\Omega)$ such that $\mathcal{A}(\gamma)\leq\mathcal{A}_{\rm Hopf}(X)$. In particular, (2.4) holds.
The inequality $c_{\rm Hopf}(X)\leq\mathcal{A}_{\rm Hopf}(X)$ is a direct consequence of the fact that (2.4) holds for every dynamically convex star-shaped domain $\Omega$ that symplectically embeds into $X$. The inequality $\mathcal{A}_{\rm Hopf}(X)\leq c_{\rm Hopf}(X)$ follows from the definition of $c_{\rm Hopf}$. The proof of Step 2 is complete.
@@
There are two simple Reeb orbits in $\partial E(a,b)$, given by the circles $(z_{2}=0)$ and $(z_{1}=0)$. These Reeb orbits have actions $a$ and $b$ respectively, and they are both Hopf orbits. If $a/b$ is rational then there are additional simple Reeb orbits, but these do not have minimal action. Using Step 2, we conclude that
@@
To prove that $c_{\operatorname{Hopf}}(Z(R))=R$, let $\Omega\subset\mathbb{R}^{4}$ be a dynamically convex domain and suppose that there exists a symplectic embedding $\varphi:\Omega\hookrightarrow Z(R)$. We need to show that $\mathcal{A}_{\operatorname{Hopf}}(\Omega)\leq R$. Given $\varepsilon>0$, there exists $C>>0$ such that $\varphi(\Omega)\subset E((1+\varepsilon)R,C)$. By Step 2, Monotonicity, and equation (2.5), it follows that
@@
Step 4
To complete the proof of Theorem 1.9, we need to show that if $X\subset\mathbb{R}^{4}$ is a dynamically convex star-shaped domain, then there exists $\gamma\in\mathcal{P}_{\rm Hopf}(\partial X)$ with $\operatorname{CZ}(\gamma)=3$ such that $\mathcal{A}(\gamma)=\mathcal{A}_{\rm Hopf}(X)$.
Let $\{\gamma_{n}\}_{n=1,\ldots}$ be a sequence in $\mathcal{P}_{\rm Hopf}(\partial X)$ satisfying $\mathcal{A}(\gamma_{n})<e\mathcal{A}_{\operatorname{Hopf}}(X)\ \forall n$ and $\lim_{n\to\infty}\mathcal{A}(\gamma_{n})=\mathcal{A}_{\operatorname{Hopf}}(X)$. We remark that by a compactness argument and the Arzelà-Ascoli theorem, one can pass to a subsequence so that $\gamma_{n}$ converges to a Reeb orbit with action $\mathcal{A}_{\operatorname{Hopf}}(X)$. However, this not enough to conclude the proof of the theorem, because the limiting Reeb orbit might not be simple or have Conley-Zehnder index $3$. We will instead need to take a limit of a sequence of Reeb orbits of different contact forms.
Let $\lambda$ denote the contact form on $\partial X$. We can construct a sequence of functions $f_{n}:\partial X\to(0,1)$ with $f_{n}\to 0$ in $C^{\infty}$ such that:
@@
  • The Reeb vector field of $e^{f_{n}}\lambda$ is tangent to $\gamma_{n}$, so that the loop $\gamma_{n}$, suitably reparametrized, defines a periodic orbit of the Reeb flow of $e^{f_{n}}\lambda$.
The existence of $f_{n}$ is elementary and somewhat standard, see [convex, Section 6]. Specifically, Lemma 6.8 from [convex] contains this existence statement where the corresponding orbit is, in addition, assumed to have ${\rm CZ}=3$, but the perturbation argument is independent of this latter assumption. The idea is quite simple. If $\hat{g}_{n}:\partial X\to\mathbb{R}$ satisfies $\hat{g}_{n}|_{\gamma_{n}}=1$ and $d\hat{g}_{n}|_{\gamma_{n}}=0$, then $\gamma_{n}$ persists as a periodic Reeb orbit of $\hat{g}_{n}\lambda$. Moreover, $\hat{g}_{n}$ can be chosen close to $1$ in $C^{\infty}$, and so that its second derivatives along $\gamma_{n}$ perturb the linearized dynamics and make $\gamma_{n}$, together with all its iterates, nondegenerate periodic orbits. Then, by the usual result, one can $C^{\infty}$-slightly perturb $\hat{g}_{n}$ to $\tilde{g}_{n}$ so that $\tilde{g}_{n}\lambda$ is a nondegenerate contact form. Since $\gamma_{n}$ is a nondegenerate periodic Reeb orbit of $\hat{g}_{n}\lambda$, it gets slightly perturbed to a periodic Reeb orbit of $\tilde{g}_{n}\lambda$ which can, via a $C^{\infty}$-small contact isotopy, be brought back to $\gamma_{n}$. After this last transformation, the contact form assumes the form $g_{n}\lambda$ for some $g_{n}$ $C^{\infty}$-close to $1$. Take $\Delta_{n}>0$ small so that $g_{n}+\Delta_{n}>1$ and define $f_{n}$ by $e^{f_{n}}=g_{n}+\Delta_{n}$.
The contact forms $e^{f_{n}}\lambda$ on $X$ correspond to star-shaped domains $X_{n}$ such that $X\subset{\rm int}(X_{n})$ and $\partial X_{n}\to\partial X$ in $C^{\infty}$ as hypersurfaces. Note that
@@
(2.7) $\lim_{n\to\infty}\mathcal{A}(\widetilde{\gamma}_{n})=\mathcal{A}_{\operatorname{Hopf}}(X)$
since $\max(f_{n})\to 0$.
Now for each $n$, choose a constant $\epsilon_{n}\in(0,\min(f_{n}))$. The contact form $e^{f_{n}-\epsilon_{n}}\lambda$ is nondegenerate and dynamically convex up to action $e^{2}\mathcal{A}_{\operatorname{Hopf}}(X)$, since it is a constant scaling of $e^{f_{n}}\lambda$. This contact form corresponds to a rescaling $X^{\prime}_{n}$ of $X_{n}$ satisfying $X\subset{\rm int}(X^{\prime}_{n})$, $X^{\prime}_{n}\subset{\rm int}(X_{n})$, and $\partial X^{\prime}_{n}\to\partial X$ in $C^{\infty}$. By Lemma 2.1 and (2.6), there exists a suitable Hopf orbit $\gamma_{n}^{\prime}\in\mathcal{P}_{\operatorname{Hopf}}(\partial X^{\prime}_{n})$ such that
(2.8) $\mathcal{A}(\gamma_{n}^{\prime})\leq\mathcal{A}(\widetilde{\gamma}_{n})$
and $\operatorname{CZ}(\gamma_{n}^{\prime})=3$. An application of Lemma 2.2 now gives the desired periodic orbit $\gamma\in\mathcal{P}_{\rm Hopf}(\partial X)$ satisfying ${\rm CZ}(\gamma)=3$ and $\mathcal{A}(\gamma)=\mathcal{A}_{\rm Hopf}(X)$. ∎
We now prepare for the proof of Lemma 2.1, which occupies the rest of Section 2.
@@ 2.2. Reeb orbits, asymptotic operators, and Conley-Zehnder indices
The following are some key facts about the spectral theory of the asymptotic operator, which follow from perturbation theory as explained in [props2, §3]. The asymptotic operator has discrete spectrum, consisting of eigenvalues whose geometric and algebraic multiplicities coincide, and the spectrum accumulates at $\pm\infty$. The eigenvectors are nowhere vanishing sections of $\gamma(T\cdot)^{*}\xi$ since they solve linear ODEs. Hence they have well-defined winding numbers with respect to a $d\lambda$-symplectic trivialization $\tau$ of $\gamma(T\cdot)^{*}\xi$. The winding number of an eigenvector depends only on the eigenvalue. So to an eigenvalue $\nu$ one can associate the winding number
@@
(2.9) $\operatorname{CZ}^{\delta}_{\tau}(\gamma)=2\alpha^{<\delta}_{\tau}(\gamma)+p^{\delta}(\gamma).$
We note that $\operatorname{CZ}_{\tau}^{\delta}$ is a lower semi-continuous function of the asymptotic operator $L_{\gamma,J}$ with respect to the $C^{\infty}$ norm. We define the Conley-Zehnder index $\operatorname{CZ}_{\tau}(\gamma)=\operatorname{CZ}_{\tau}^{0}(\gamma)$, i.e. by taking $\delta=0$ in (2.9).
@@
$\operatorname{CZ}(\gamma)=\operatorname{CZ}_{\tau}^{0}(\gamma).$